OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 12 cm

OpenStudy (anonymous):

a^2 + b^2 = c^2 Since it's isosceles a=b a^2 + a^2 = 12^2 2a^2 = 144 a^2 =72 a=8.49

OpenStudy (anonymous):

yea.so a=b=sqr(72)=8.49...........jst apply pythagoras.theorm